# Direct Preference Knowledge Distillation for Large Language Models

### v1

@@ Abstract
In the field of large language models (LLMs), Knowledge Distillation (KD) is a critical technique for transferring capabilities from teacher models to student models. However, existing KD methods face limitations and challenges in distillation of LLMs, including efficiency and insufficient measurement capabilities of traditional KL divergence. It is shown that LLMs can serve as an implicit reward function, which we define as a supplement to KL divergence. In this work, we propose Direct Preference Knowledge Distillation (DPKD) for LLMs. DPKD utilizes distribution divergence to represent the preference loss and implicit reward function. We re-formulate KD of LLMs into two stages: first optimizing and objective consisting of implicit reward and reverse KL divergence and then improving the preference probability of teacher outputs over student outputs. We conducted experiments and analysis on various datasets with LLM parameters ranging from 120M to 13B and demonstrate the broad applicability and effectiveness of our DPKD approach. Meanwhile, we prove the value and effectiveness of the introduced implicit reward and output preference in KD through experiments and theoretical analysis. The DPKD method outperforms the baseline method in both output response precision and exact match percentage. Code and data are available at https://aka.ms/dpkd.
@@ 1 Introduction
In the era of Large language models (LLMs), a series of models and techniques have demonstrated great capabilities Zhang et al. (2022a); Chowdhery et al. (2023) , and their powerful capabilities are usually ascribed to the increase in the size of training data and model scale Kaplan et al. (2020); Anil et al. (2023). However, the increase in the size of large models also brings expensive computing costs and difficulties Hoffmann et al. (2022). How to maintain performance and conduct efficient training while increasing the scale of language models has become an important issue. Knowledge distillation (KD; Hinton et al., 2015) trains a relatively compact student model by simulating the output distribution and behavior of the teacher model, which is an effective method under limited computing resources. For white-box models, we can have access to richer information such as the distribution of logits of the model Gou et al. (2021). This provides us with more value because the distribution of the hidden layer represent the underlying knowledge of LLMs, leading to better effect and output in knowledge transfer Wang et al. (2020).
@@
In this work, we propose Direct Preference Knowledge Distillation (DPKD) for knowledge distillation of LLMs. It is shown that LLMs can serve as an implicit reward function Yuan et al. (2024); Rafailov et al. (2024b). Due to the deficiency of KL divergence Gu et al. (2023); Huang et al. (2022), we define implicit reward function as a supplement to the KL distance. We re-formulate the KD of white-box LLMs as follows, first maximize the optimization function consisting of implicit reward and reverse KL divergence, and then improve the preference probability of teacher outputs over student outputs. These settings compensate for the shortcomings of KL divergence while preserving high training efficiency. We derive the training objective to obtain a concise final form, and prove the role of the reward and preference form we introduced in KD through theoretical derivation.
Finally, we conducted experiments on our DPKD method in the instruction tuning topic. We use different families of LLMs including GPT-2 Radford et al. (2019) and OPT Zhang et al. (2022b) with parameter sizes ranging from 120M to 13B and multiple datasets to validate our method. We evaluate the results with RougeL Lin (2004) metric , which is commonly used in instruction fine-tuning tasks. Our results show that the DPKD method outperforms the baseline method and has advantages over a wide range of generation lengths. We also conducted experiments on the implicit reward function in the reformulation of distillation to verify effectiveness, and experiments on the forms of preference to show the potential of preference modeling in KD and provide ideas for subsequent research. Our contributions are thus three-fold:
@@ 2.1.2 Preference Model
$=\sigma\left[r\left(y_{1}\right)-r\left(y_{2}\right)\right],$
@@ 2.2 DPKD: Direct Preference Knowledge Distillation
In the context of knowledge distillation of LLMs, the distribution difference measured by KL divergence is usually considered to be the only criterion for the distillation target. But KL divergence is not enough. From the perspective of measuring distributions, KL divergence is 0 if and only if $q_{\theta}$ and $p$ are the same distribution in the case of discrete variables, or almost identical in the case of continuous variables. KL divergence measures the difference between two distributions, but it is not really a distance because it is asymmetric Manning and Schutze (1999) . Therefore, some researchers have improved the KL divergence by combining two symmetrical items (Jensen-Shannon divergence Nielsen (2021)). In the field of KD, researchers have explored the inadequacy of KL divergence in KD Cho and Hariharan (2019); Mirzadeh et al. (2020). Some works Gu et al. (2023); Wu et al. (2024) hope to redesign KL divergence, and some worksHuang et al. (2022) add new objectives besides KL divergence.
@@
where $q^{*}(y|x)\mathrel{\coloneqq}\frac{1}{Z(x)}p(y|x)\exp\left(\frac{1}{\beta}r(x,y)\right)$, and $Z\left(x\right)$ is the normalized function of the distribution, which is required to be independent of $\theta$ and $y$. It is defined as $Z(x)\mathrel{\coloneqq}\sum_{y}p(y|x)\exp\left(\frac{1}{\beta}r(x,y)\right)$. Following the work, we can derive the optimal solution of Equation 3:
@@ 2.2.1 Learning Objective
$\log\sigma\left(\frac{\beta}{\lvert y_{t}\rvert}\log\frac{q_{\theta}\left(y_{t}\right)}{p\left(y_{t}\right)}-\frac{\beta}{\lvert y_{s}\rvert}\log\frac{q_{\theta}\left(y_{s}\right)}{p\left(y_{s}\right)}\right).$ (9)
@@
0:    Instruction tuning datasets $D$; Pre-training corpus $D_{p}$; Fine-tuned teacher model with distribution $p$;Initialized student model conducted SFT on $D_{p}$ with outpu distribution $q_{\theta}$;
@@
4:        Computes four log items $\log p(y_{t})$, $\log p(y_{s})$, $\log q_{\theta}(y_{t})$ and $\log q_{\theta}(y_{s})$
5:        Compute distill loss $\mathcal{L}_{kd}=-\sum_{\left(x\sim\mathcal{D}\right)}$ $\log\sigma\left(\frac{\beta}{\lvert y_{t}\rvert}\log\frac{q_{\theta}\left(y_{t}\right)}{p\left(y_{t}\right)}-\frac{\beta}{\lvert y_{s}\rvert}\log\frac{q_{\theta}\left(y_{s}\right)}{p\left(y_{s}\right)}\right)$
6:        Compute language modeling loss $\mathcal{L}_{pt}=-\sum_{d\sim\mathcal{D}_{p}}\log q_{\theta}(d)$
7:        Compute loss $\mathcal{L}=\mathcal{L}_{kd}+\lambda\cdot\mathcal{L}_{pt}$ and gradient $\nabla_{\theta}\mathcal{L}$
8:        Gradient Update $\theta^{(t+1)}=\theta^{(t)}-\alpha\nabla_{\theta}\mathcal{L}$
@@
11:  return  Student model parameter $\theta$;
@@ 3.2 Optimizing DPKD Is Optimizing The Q Function
From another perspective of Markov Decision Process (MDP), the process of sequence generation by the model is regarded as the generation process of a Markov decision chain. The state is denoted as $s_{t}=\{x_{0},\ldots,x_{l-1},y_{0},\ldots,y_{t-1}\}$, and the action is $y_{t}\in\mathbf{V}$, which is chosen based on generated sequence $s_{t}$. The transition function from the current state to the next state is the LLMs we selected.
We analyze the implicit reward function of DPKD from $Q$ function, which is a perspective in reinforcement learning. $Q$ function is specifically explained as follows: $Q\left(s,a\right)$ refers to the expectation of the benefit obtained by taking action $a$ (a word selected from the vocabulary, obtained by taking the maximum from the output distribution or sampling by top-k) in state $s$ (generated sequence $\{x,y\}$). The general represent of $Q$ is:
@@
where $V^{*}(\mathbf{s})$ is defined to be zero if $s$ is the end of sequence (EOS). Following the work , we can use the $Q$ function instead of $r$ and substitute it into the Bellman equation.
Now we can use the $Q$ function to re-derive the DPKD method. We use the $Q$ function and the $V$ function to represent the sum of the reward function written at the current time step t. By transforming the Equation 14 and summing over time t, and substitute Equation 13 into the resulting sum, we could obtain:
@@
where $\tau$ refers to the sequence trajectory generated by the model, and in the sequence generation task of the large model, it refers to the generated text. Replace $\tau^{t}$ and $\tau^{s}$ in the formula with $y_{t}$ and $y_{s}$ respectively, and the above formula is exactly the same as Equation 6. Thus we get the complete process of deriving the DPKD of this work from the $Q$ function. Therefore, we can link the DPKD method of this work with the concept of $Q$ function in reinforcement learning and Markov decision chain.
@@ 3.3 Your Language Model Is Secretly Implicit Reward Function During KD
From the experimental results, we implicitly optimized a reward function in the KD process by redefining the formula and objectives of KD, and our experimental results show the role of the reward function in the KD training process. This shows that the KD in our experiment is actually optimizing a reward function, which is an unobserved perspective.
From the perspective of theoretical analysis, the reward function not only serves as a weight term in the gradient of our DPKD, but also as a hub connecting our DPKD method and the Q function of reinforcement learning.
Furthermore, from the results in Section 4.4, we can see that experiments with different preference expressions related to reward functions show quite different results, and some preference expressions even show promising results, even though we did not do much parameter fine-tuning. This shows that using reward functions to express preferences in the knowledge distillation process is effective, and there is still a lot of value to explore, which also provides ideas for subsequent work.
@@ 4.1 Experiments Setup
We conduct experiments and analysis on the task of instruction tuning. The topics of instruction tuning range from summarizing to completing the request, and the model is required to complete the response according to instruction, prompt, and input. We specifically select data with instructions and outputs longer than 10 words in the dataset to ensure that the model aims to generate high-quality results that are long enough. The RougeL Lin (2004) score is used to measure the quality of model generation, as Wang et al. (2022) has shown that RougeL is suitable for large-scale instruction tuning evaluation.
@@ 4.2 Results
We conducted experiments based on the Algorithm 1 and the experimental settings in Section 4.1, and the experimental results are shown in Table 1. In this section, we conduct a preliminary analysis of the results including the RougeL score, the GPT-4 score, and an example of the implicit reward function introduced in this article.
From the experimental results we could see that our method shows great performance compared with the various baseline method. We conduct KD experiments on datasets with different distributions and LLMs of different sizes, showing that our method is applicable to a wide range of models and data and retains good performance. In some specific datasets, the student model trained by our knowledge distillation method is even close to the performance of the teacher model.
@@
We collected the output texts of DPKD and the baseline method of GPT-2 Base, and scored them together with the labels of the dataset by GPT-4 to evaluate the quality of the instruction tracking generated text of each method compared with the label. Due to the input context text restrictions of GPT-4, we used the generation results and labels of the first 100 datasets of DollyEval. Although the output results of each method are still a certain distance away from the label, the generation results of the method proposed in this paper are close to the evaluation of the label, which is higher than baselines.
@@ 4.3 Reward Observation
We analyzes the implicit reward function involved in this paper, and analyzes the role of the reward function introduced in KD from both theoretical and experimental perspectives. According to the derivation of the gradient in Section 3.1, the reward function plays the role of a weight in the gradient update process. When the model parameter’s estimate of the reward function indicates that the current output is biased towards error, the value of the weight will be larger, and thus the gradient will be larger. When the current performance of the model is already better and the output distributions of the student model and the teacher model are close, the weight term we mentioned earlier will be low, so the model will tend to converge.
We hope to examine the trend of KL divergence and reward function of DPKD. We use an unbiased approximation to the reward function, which is $\hat{r}^{*}(x,y)=\beta\log\frac{q_{\theta}(y|x)}{p(y|x)}$ Based on this and the definition of reward estimation, we plot the two variables of reward function and KL divergence for the same model (here we use the GPT-2 Base model) at different training periods. We can get the following results.
@@
We visualize the trend of reward during training and the reverse KL divergence distance between the student and teacher models over training epochs in Figure 2. From the visualization results, we can see that although some model checkpoints are very close to the KL divergence of the teacher model (all within the range of $0.04\pm 0.02$), the difference in the reward function value still causes the difference in RougeL scores. From the curve trend in the figure, it can be seen that the optimal point of model training is taken at a left-point arc vertex, that is, the student model has the lowest KL divergence from the teacher and the reward function is at a higher value. The size of the reward function may be related to the hyperparameter $\beta$ in training, which is used to adjust the relative proportion of the KL divergence and the reward function.
Figure 3 shows the changes in the implicit reward function, KLD, and rKLD during training. We can see that the estimated implicit reward gradually increases within a range. The trends of KLD and rKLD are roughly the same, but the specific values are slightly different, which is consistent with the conclusions of previous work Gu et al. (2023); Wu et al. (2024).

### v2

@@ Abstract
In the field of large language models (LLMs), Knowledge Distillation (KD) is a critical technique for transferring capabilities from teacher models to student models. However, existing KD methods face limitations and challenges in the distillation of LLMs, including efficiency and insufficient measurement capabilities of traditional Kullback-Leibler (KL) divergence. It is shown that LLMs can serve as an implicit reward function, which we define as a supplement to KL divergence. In this work, we propose Direct Preference Knowledge Distillation (DPKD) for LLMs. DPKD utilizes distribution divergence to represent the preference loss and the implicit reward function. We re-formulate KD of LLMs into two stages: first optimizing an objective consisting of implicit reward and reverse KL divergence and then improving the preference probability of teacher outputs over student outputs. We conducted experiments on various datasets with LLM parameters ranging from 120M to 13B and demonstrate the broad applicability and effectiveness of our DPKD approach. Meanwhile, we prove the value and effectiveness of the introduced implicit reward and output preference in KD through experiments and theoretical analysis. The DPKD method outperforms the baseline method in both output response precision and exact match percentage.
@@ 1 Introduction
In the era of Large Language Models (LLMs), a series of models and techniques have demonstrated great capabilities Zhang et al. (2022a); Chowdhery et al. (2023), and their powerful capabilities are usually ascribed to the increase in the size of the training data and the scale Kaplan et al. (2020); Anil et al. (2023). However, the increase in the size of large models also brings expensive computing costs and difficulties Hoffmann et al. (2022). How to maintain performance and conduct efficient training while increasing the scale of language models has become an important issue. Knowledge distillation (KD; Hinton et al., 2015) trains a relatively compact student model by simulating the output distribution and behavior of the teacher model, which is an effective method under limited computing resources.
@@
In this work, we propose Direct Preference Knowledge Distillation (DPKD) for the knowledge distillation of LLMs. It is shown that LLMs can serve as an implicit reward function Yuan et al. (2024); Rafailov et al. (2024b). Due to the deficiency of KL divergence Gu et al. (2023); Huang et al. (2022), we define implicit reward function as a supplement to the KL distance. We re-formulate the KD of white-box LLMs as follows, first maximize the optimization function consisting of implicit reward and reverse KL divergence, and then improve the preference probability of teacher outputs over student outputs. These settings compensate for the shortcomings of KL divergence while preserving high training efficiency. We derive the training objective to obtain a concise final form, and demonstrate the significance of the reward and preference form we introduced in KD through theoretical derivation.
Finally, we conducted experiments on our DPKD method in the instruction tuning task. We utilized various families of LLMs including GPT-2 Radford et al. (2019) and OPT Zhang et al. (2022b) with parameter sizes ranging from 120M to 13B and multiple datasets to validate our method. We evaluate the results with Rouge-L Lin (2004) metric. Our results show that the DPKD method outperforms the baseline and has advantages over a wide range of generation lengths. Additionally, we conducted experiments on the implicit reward function in the reformulation of distillation to verify its effectiveness, and experiments on the forms of preference to show the potential of preference modeling in KD and provide ideas for subsequent research. Our contributions are thus three-fold:
@@ 2.1.2 Preference Model
$=\sigma\left(r\left(y_{1}\right)-r\left(y_{2}\right)\right),$
@@ 2.2 DPKD: Direct Preference Knowledge Distillation
In the context of knowledge distillation of LLMs, the difference in distribution measured by KL divergence is often regarded as the only criterion for the distillation target. But KL divergence is insufficient. KL divergence measures the difference between two distributions, but it is not really a distance because it is asymmetric Manning and Schutze (1999). Researchers have improved the KL divergence by combining two symmetrical items (Jensen-Shannon divergence Nielsen (2021)). In the field of KD, researchers have explored the inadequacy of KL divergence in KD Cho and Hariharan (2019); Mirzadeh et al. (2020). Some works Gu et al. (2023); Wu et al. (2024) hope to redesign KL divergence, and some worksHuang et al. (2022) add new objectives besides KL divergence. We conducted a toy experiment to further illustrate the inadequacy of KLD in Figure 2 and Section 4.3.
@@
where $q^{*}(y|x)\mathrel{\coloneqq}\frac{1}{Z(x)}p(y|x)\exp\left(\frac{1}{\beta}r(x,y)\right)$, and $Z\left(x\right)$ is the scaling function of the distribution, which is required to be independent of $\theta$ and $y$. It is defined as $Z(x)\mathrel{\coloneqq}\sum_{y}p(y|x)\exp\left(\frac{1}{\beta}r(x,y)\right)$. Following the work, we can derive the optimal solution of Equation 3:
@@
Given the same prompt x, the outputs of the student model and the teacher model are denoted as $y_{s}$ and $y_{t}$ respectively. The purpose of KD is to fit the distribution of the student model to teacher model, which can also be understood as we expect the student model to have a greater probability of outputting results similar to those of the teacher model. From the BT model, we can obtain the following probability:
@@ 2.2.1 Learning Objective
$\mathcal{L}=-\mathbb{E}\left[\log\sigma\left(\frac{\beta}{\lvert y_{t}\rvert}\log\frac{q_{\theta}\left(y_{t}\right)}{p\left(y_{t}\right)}-\frac{\beta}{\lvert y_{s}\rvert}\log\frac{q_{\theta}\left(y_{s}\right)}{p\left(y_{s}\right)}\right)\right].$ (9)
@@
0:    Instruction tuning datasets $D$; Pre-training corpus $D_{p}$; Fine-tuned teacher model with distribution $p$;Initialized student model conducted SFT on $D_{p}$ with output distribution $q_{\theta}$;
@@
4:        Compute four log items $\log p(y_{t})$, $\log p(y_{s})$, $\log q_{\theta}(y_{t})$ and $\log q_{\theta}(y_{s})$;
5:        Compute distill loss $\mathcal{L}_{kd}=-\sum_{\left(x\sim\mathcal{D}\right)}$; $\log\sigma\left(\frac{\beta}{\lvert y_{t}\rvert}\log\frac{q_{\theta}\left(y_{t}\right)}{p\left(y_{t}\right)}-\frac{\beta}{\lvert y_{s}\rvert}\log\frac{q_{\theta}\left(y_{s}\right)}{p\left(y_{s}\right)}\right)$;
6:        Compute language modeling loss $\mathcal{L}_{pt}=-\sum_{d\sim\mathcal{D}_{p}}\log q_{\theta}(d)$;
7:        Compute loss $\mathcal{L}=\mathcal{L}_{kd}+\lambda\cdot\mathcal{L}_{pt}$ and gradient $\nabla_{\theta}\mathcal{L}$;
8:        Gradient Update $\theta^{(t+1)}=\theta^{(t)}-\alpha\nabla_{\theta}\mathcal{L}$;
@@
11:  return  Student model parameter $\theta$.
@@ 3.2 Optimizing DPKD Is Optimizing The Q Function
From another perspective of Markov Decision Process (MDP), the process of sequence generation by the model is regarded as the generation process of a Markov decision chain. The state is denoted as $s_{t}=\{x_{0},\ldots,x_{l-1},y_{0},\ldots,y_{t-1}\}$, and the action is $y_{t}\in\mathbf{V}$, which is chosen based on generated sequence $s_{t}$. The transition function from the current state to the next state is the LLMs we selected. We analyze the implicit reward function of DPKD from $Q$ function, which is a perspective in reinforcement learning. The general represent of $Q$ is:
@@
where $V^{*}(\mathbf{s})$ is defined to be zero if $s$ is the end of sequence (EOS). Following work Rafailov et al. (2024a), we can use the $Q$ function instead of $r$ and substitute it into the Bellman equation.
Now we can use the $Q$ function to re-derive the DPKD method. By transforming the Equation 14 and summing over time t, and substitute Equation 13 into the result, we could obtain:
@@
where $\tau$ refers to the sequence trajectory generated by the model, and in the sequence generation task of the large model, it refers to the generated text. Replace $\tau^{t}$ and $\tau^{s}$ with $y_{t}$ and $y_{s}$ respectively, and the above formula is exactly the same as Equation 6. Thus we get the complete process of deriving the DPKD of this work from the $Q$ function. Therefore, we can link the DPKD method of this work with the concept of $Q$ function in reinforcement learning and Markov decision chain.
@@ 3.3 Your Language Model Is Secretly Implicit Reward Function During KD
We implicitly optimized a reward function during the KD process by redefining the formula and objectives of KD, and our experimental results demonstrate the significance of the reward function in the KD training process. From the perspective of theoretical analysis, the reward function serves not only as a weight term in the gradient of our DPKD but also as a crucial link between our DPKD method and the Q function in reinforcement learning.
Furthermore, the results presented in Section 4.4 indicate that various preference expressions related to reward functions yield significantly different results. Notably, some preference expressions demonstrate promising results, despite minimal parameter fine-tuning. This suggests that employing reward functions to articulate preferences in the knowledge distillation process is effective and highlights the potential for further exploration, which also offers valuable insights for future research.
@@ 4.1 Experiments Setup
We conduct experiments and analyses on the task of instruction tuning. The task of instruction tuning range from summarizing to completing requests, requiring the model to generate responses based on the provided instructions, prompts, and inputs. To ensure that the model produces high-quality results of sufficient length, we specifically select data with instructions and outputs exceeding 10 words in the dataset. The Rouge-L score Lin (2004) is employed to assess the quality of model generation, as Wang et al. (2022) has shown that Rouge-L is appropriate for large-scale instruction tuning evaluation.
@@ 4.2 Results
We conducted experiments based on Algorithm 1 and the experimental settings in Section 4.1, and the experimental results are presented in Table 1. In this section, we conduct a preliminary analysis of the results including the RougeL score, the GPT-4 score, and an example of the implicit reward function introduced in this article.
From the experimental results, we could see that our method demonstrates great performance compared to the various baseline. We conduct KD experiments on datasets with different distributions and LLMs of different sizes, showing that our method is applicable to a wide range of models and data and retains good performance. In some specific datasets, the student model trained by our knowledge distillation method is even close to the performance of the teacher model.
@@
We collected the output texts of DPKD and the baseline method of GPT-2 Base, and evaluated them together with the labels of the dataset by GPT-4 to evaluate the quality of the instruction tracking generated text of each method in Figure 1. Although the output results of each method are still a certain distance away from the label, the generation results of the method proposed in this paper are closer to the label evaluations and surpass those of the baselines.
@@ 4.3 Reward Observation
In this section, our analysis highlights the inadequacies of rKLD and emphasizes the significance of the reward function introduced in this paper. We conducted an experiment to further illustrate the deficiencies of rKLD discussed in Section 2.2 and the corresponding changes in reward and Rouge-L. As shown in Figure 2, we added random noise to the basic model and simultaneously calculated both the rKLD and the estimated implicit reward presented in this paper during inference. Notably, while the rKLD fluctuates at a relatively low value of 0.001, the Rouge-L score exhibits a fluctuation as high as 6.23.
Figure 2: Illustration of the relation of rKLD, implicit reward and Rouge-L. We construct differentiated models by adding random noise to the base model. Lighter colors indicate higher Rouge-L scores.
@@
Furthermore, we analyze the implicit reward function discussed in this paper and examine the role of the reward function introduced in KD from both theoretical and experimental perspectives. According to the gradient derivation presented in Section 3.1, the reward function acts as a weight in the gradient update process. When the estimate of the reward function indicates that the current output is biased towards an error, the weight value will increase, resulting in a larger gradient. Conversely, when the model’s performance improves and the output distributions of the student and teacher models become similar, the previously mentioned weight term will decrease, leading the model to converge.
We aim to analyze the trend of KL divergence and the reward function of DPKD. To achieve this, we utilize an unbiased approximation of the reward function, which is $\hat{r}^{*}(x,y)=\beta\log\frac{q_{\theta}(y|x)}{p(y|x)}$. Based on this and the definition of reward estimation, we plot the reward function and KL divergence for the same model (in this case, we use the GPT-2 Base) at different training periods.
We illustrate the trend of reward during training and the reverse KL divergence distance between the student and teacher models across training epochs in Figure 3. From the visualization results, we observe that while some model checkpoints exhibit very close KL divergence values to those of the teacher model (all within the range of $0.04\pm 0.02$), the differences in reward values still lead to variations in the RougeL scores. From the curve trend in the figure, it can be seen that the optimal point of model training occurs at a left-point arc vertex, while the student model has the lowest KL divergence from the teacher and the reward function is at a higher value. The magnitude of the reward function may be related to the hyperparameter $\beta$ used during training, which adjusts the relative proportions of the KL divergence and the reward function.
Figure 4 illustrates the changes in the implicit reward function, KLD, and rKLD during training. We can see that the estimated implicit reward gradually increases within a specific range. The trends of KLD and rKLD are generally similar, but the exact values differ slightly, which is consistent with the conclusions of previous work Gu et al. (2023); Wu et al. (2024).